OpenStudy (anonymous):

OMG CAN SOMEONE HELP ME, NOW???????????????//////// DO YOU GUYS KNOW TRIG? :( FOR THE THIRD TIME: The expression 3N + 2/ (N-1)(N-2) can be written in the following form: A/(N-1)-B/(N-2) Compute A^2 + B ^2

OpenStudy (anonymous):

We don't have to help. Keep in mind we might not know it either.

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

I'm going to keep asking until I find someone who knows. :)

OpenStudy (aum):

\[ \frac{ 3N + 2}{ (N-1)(N-2) } = \frac{A}{N-1} - \frac{B}{N-2} = \frac{A(N-2)-B(N-1)}{(N-1)(N-2)} = \\ \frac{ 3N + 2}{ (N-1)(N-2) } = \frac{(A-B)N+(B-2A)}{(N-1)(N-2)} \\ A- B = 3 \\ B-2A = 2 \\ \text{Solve for A and B} \]

OpenStudy (jdoe0001):

hmmm that doesn't look like trig btw

OpenStudy (anonymous):

Then what is it? :|

OpenStudy (anonymous):

It's not trig, but I am taking trig and that's what my teacher gave as hmw lol

OpenStudy (aum):

What did you get for A and B?

OpenStudy (aum):

A - B = 3 -2A + B = 2 Add them and solve for A. Then sub A in first equation and solve for B.

OpenStudy (anonymous):

B = -8 A = 11 ?

OpenStudy (jdoe0001):

hmmm I don't see it coming out.. myself

OpenStudy (aum):

A = -5 B = -8